OpenStudy (anonymous):

can you show me how to solve 2x-y=3 4x+5y=6 using the substitution method. I am stuck on this question for my math hw. pretty please!

OpenStudy (anonymous):

1. Solve for a variable for ONE of the equations - Get x, or y, by its self 2. Plug in the equation you found for the variable of the other equation & solve. You will get a value for the variable. 3. Put this back into the equation you found in step 1 and get the answer for the other variable. Example: x+y = 12 y - 6 = 2x 1. x = 12-y 2. y - 6 = 2(12-y) y - 6 = 24 - 2y 3y - 6 = 24 3y = 30 y = 10 3. x= 12- 10 x = 2 Answer: x = 2, y = 10. Now try and use these steps for your problem. :)